### v2

@@ 4 Experiments
In the following, we first introduce the datasets in Sections 4.1, then describe the experimental methodology for both registered and unregistered settings in Section 4.2. Finally, we present a comprehensive evaluation in Section 4.3.
4.1 Datasets
We use three standard datasets for training and quantitative evaluation in 3D speech-driven motion synthesis: VOCAset [10], BIWI [15], and Multiface [56].
VOCAset consists of mesh sequences from 12 actors performing 40 speeches, recorded at 60 fps. Each sequence lasts approximately 3 to 5 seconds, with meshes in FLAME [24] topology, which has 5,023 vertices and 9,976 faces.
BIWI includes data from 14 subjects who articulate 40 sentences each, sampled at 25 fps. Each sentence is about 5 seconds long, and the meshes are registered to a fixed topology. Due to GPU limitations, we utilize a downsampled version of this dataset, denoted as BIWI6, which features a fixed topology with 3,895 vertices and 7,539 faces.
Multiface contains data from 13 identities performing up to 50 speeches of approximately 4 seconds each, recorded at 30 fps. The original Multiface meshes are registered to a fixed topology with 5,471 vertices and 10,837 faces.
For training, the meshes from both Multiface and BIWI6 are scaled and aligned to match the VOCAset meshes. To facilitate a fair comparison with existing methods, we train four distinct models: one using all the three datasets (multi-dataset), and three using each dataset separately (single-dataset).
4.2 Experiments methodology
@@
Registered setting: in this setting, all meshes inside a given sequence are aligned to a common topology. This allows using a supervised loss such as the $\mathcal{L}_{MSE}$, $\mathcal{L}_{M}$, $\mathcal{L}_{Vel}$ and $\mathcal{L}_{Cos}$. Even though the topology may vary from one sequence to another, when training a single model on several registered datasets the only requirement is to rigidly align and scale the data with a reference. For our experiments, we rigidly aligned all the meshes with the VOCAset meshes.
Unregistered setting. This setting is similar to the registered setting, however we replace $\mathcal{L}_{MSE}$ with the dynamic chamfer distance $\mathcal{L}_{CD}$.
Figure 3: Original VOCAset mesh (left) and remeshed version (right) for unregistered training.
To showcase the results of a training procedure on unregistered meshes, we chose to randomly remesh VOCAset by upsampling and downsampling the meshes as depicted in Figure 3. In doing so, each mesh of each sequence has its own structure, different from all the others.
@@
For other settings, such as training directly on scans, one should take care about rigid alignment with the training data and the existence of a clear aperture for the mouth. Also, downsampling the training meshes is a pre-processing step that can make training faster.
4.3 Experimental Results
4.3.1 Registered Training
@@
Table 1 shows the performance of ScanTalk when trained with various loss combinations across all three registered datasets in a multidataset setting. The table reports both the proposed metrics and the standard ones. We compare against state-of-the-art methods trained separately on each dataset. At a first glance, it can be easily noted that none of the compared approaches outperforms the others with respect to all the metrics. In fact, while standard metrics (LVE, MVE, FDD) mostly focus on geometric measures, they do not explicitly account for the dynamics of the animation. The proposed ones instead (DTW, DFD, $\delta_{M}$, $\delta_{Cd}$) provide such information. This points towards the idea that there is still large room for improvements, but without proper measurements such assessment is difficult.
Figure 4: Comparison of ScanTalk with state-of-the-art methods on a VOCAset test set sequence, focusing on the movement of the lips ($y$-coordinate). The $y$-coordinate most effectively captures lips movements, which show a larger extent in the vertical dimension. The additional losses used to train ScanTalk help in better following the lips dynamics.
One interesting result that supports our claim is that in Table 1 vanilla ScanTalk performs better than ScanTalk trained with the additional losses (masked, velocity and cosine) in terms of geometric measures such as LVE. However, if one looks at the charts in Figure 5, which reports the generated $y$-coordinate movements of the upper and lower lips against ground truth, can clearly observe that the dynamics of the animation is better modeled when the additional losses are used. This is reflected in lower DTW, DFD, $\delta_{M}$ and $\delta_{Cd}$ measures in Table 1, which represent a quantitative alternative to the qualitative results in the charts.
@@
Table 1: Performance metrics across different datasets. MV stands for mask+velocity, MVC stands for mask+velocity+cosine, and CH stands for chamfer. Each subtable presents the results for a specific dataset. While each state-of-the-art model is trained on a single dataset, Scantalk versions are trained in a multidataset setting.
VOCAset
LVE $\downarrow$ MVE $\downarrow$ FDD $\downarrow$ DTW DFD $\delta_{M}$ $\delta_{Cd}$
VOCA 6.993 0.983 2.662 1.773 7.413 1.394 1.019
FaceFormer 6.123 0.935 2.163 1.329 5.395 0.861 0.584
FaceDiffuser 4.350 0.901 2.437 1.734 6.831 0.811 0.606
CodeTalker 3.549 0.888 2.258 1.329 5.661 0.805 0.552
SelfTalk 5.618 0.918 2.321 1.249 5.431 0.726 0.567
ScanTalk 6.375 0.987 2.101 1.387 5.710 0.754 0.596
ScanTalk-MV 7.031 1.092 1.473 1.144 5.161 0.781 0.583
ScanTalk-MVC 7.050 1.066 1.287 1.192 5.254 0.715 0.518
ScanTalk-CH 6.572 1.004 2.305 1.529 6.236 0.771 0.587
BIWI6
LVE $\downarrow$ MVE $\downarrow$ FDD $\downarrow$ $DTW$ $DFD$ $\delta_{M}$ $\delta_{Cd}$
VOCA 5.743 2.591 41.482 1.602 7.892 1.478 0.671
FaceFormer 4.085 2.163 37.091 1.566 6.615 0.871 0.551
FaceDiffuser 4.022 2.128 39.604 1.550 6.502 0.851 0.568
CodeTalker 5.190 2.641 20.599 1.497 6.609 0.913 0.592
SelfTalk 3.628 2.062 35.470 1.598 6.932 1.031 0.659
ScanTalk 4.044 2.057 40.051 1.473 6.482 0.971 0.606
ScanTalk-MV 4.755 2.227 39.458 1.513 6.741 1.092 0.632
ScanTalk-MVC 4.448 2.096 36.653 1.469 6.712 1.016 0.619
ScanTalk-CH 4.301 2.142 34.979 1.548 6.756 0.998 0.620
Multiface
LVE $\downarrow$ MVE $\downarrow$ FDD $\downarrow$ $DTW$ $DFD$ $\delta_{M}$ $\delta_{Cd}$
VOCA 4.923 2.761 55.781 1.569 6.512 0.861 1.239
FaceFormer 2.451 1.453 20.239 0.868 4.132 0.221 0.735
FaceDiffuser 3.555 2.388 29.157 1.450 5.243 0.281 0.772
CodeTalker 4.091 2.382 47.905 1.438 5.951 0.546 0.965
SelfTalk 2.281 1.901 37.434 0.959 4.183 0.255 0.759
ScanTalk 2.435 1.678 32.202 0.900 4.081 0.233 0.746
ScanTalk-MV 2.547 1.898 15.973 0.993 4.217 0.276 0.771
ScanTalk-MVC 2.370 1.712 16.003 0.938 4.718 0.331 0.740
ScanTalk-CH 3.712 2.522 110.630 1.362 5.641 0.357 0.748
@@
Overall, Table 1 indicates that ScanTalk, trained in a multidataset setting, performs favorably with respect to state-of-the-art models, mostly on VOCAset. The newly introduced metrics though provide a more complete evaluation. For example, in VOCAset, FaceDiffuser and CodeTalker report the lowest LVE and MVE. However, their motion dynamics is quite compromised when looking at the generated animations. To support this, Figure 4 reports their motion charts for a randomly picked sequence of VOCAset. The generated motion fails to a large extent in following that of the ground-truth. Indeed, this is again reflected in worse DTW, DFD and $\delta$ measures. ScanTalk trained with the additional losses instead better captures the motion dynamics. This improvement aligns with our expectations, as the new losses were specifically designed to enhance dynamic animation realism, a nuance that the standard metrics failed to capture. This comparison highlights the limitations of standard metrics, demonstrating their inadequate performance. In contrast, the new metrics we propose offer a significantly improved assessment of lip-sync quality. Heatmaps in Figure 6(a), which show the MSE between the ground truth and predictions on VOCAset, qualitatively reveal that ScanTalk performs comparably with other state-of-the-art methods.
Figure 5: Lips motion dynamics ($y$ coordinate) of ScanTalk trained with different losses on two test set sequences for each dataset. Charts show that the mouth motion dynamics is improved by adding the velocity $\mathcal{L}_{Vel}$ and the cosine $\mathcal{L}_{Cos}$ losses.
@@
(a) Comparison of ScanTalk with state-of-the-art methods on VOCAset. Blue/red hues = lower/higher error.
@@
(b) MDS visualization of latent combinations from ScanTalk. Meshes with the same identity and audio show highly similar representations despite parameterization changes. Different identities with the same audio exhibit similar MDS trajectories
@@
In $BIWI_{6}$ and Multiface performance vary more. Specifically, ScanTalk surpasses the state-of-the-art method on $BIWI_{6}$ in DTW and DFD, but not in $\delta_{M}$ and $\delta_{Cd}$. For Multiface, ScanTalk outperforms state-of-the-art methods for the DFD metric and performs comparably in terms of other metrics. These observations align with findings from [30], and we attribute this performance discrepancy to the significant differences among the three registered datasets. VOCAset features static faces with only mouth movements, while $BIWI_{6}$ and particularly Multiface involve notable head movements in their sequences. These dataset variations contribute to the model’s reduced performance as they introduce a further level of complexity to be modeled in a unified training scheme.
Figure 5 presents a comparative analysis of ScanTalk’s predicted lip movements across two test sequences from each dataset. We observe a significant variation in accuracy among the datasets ground-truth sequences. Specifically, VOCAset exhibits the most precise lip movements compared to BIWI and Multiface. Additionally, incorporating supplementary losses during training notably enhances performance on VOCAset. However, this effect is less pronounced for BIWI and Multiface due to smaller lip motion dynamism in the sequences.
@@
Figure 6(b) presents the Multidimensional Scaling (MDS) [51] projection of three sequences of latent combinations from ScanTalk, each generated with the same audio snippet but different faces to animate. The results show that when two faces share the same identity but differ in topology, their MDS sequences are very similar, nearly overlapping. In contrast, when animating different identities with the same audio, the MDS paths exhibit similar patterns but occupy distinct positions in space. These results underscore ScanTalk capability to animate faces independently of topology, focusing on the face identity. From Table 1, it is evident that ScanTalk, when trained with the chamfer loss in a registered setting, yields results comparable to those obtained using registered losses in the same setting. This comparison underscores the robustness and adaptability of our framework.
4.3.2 Unregistered Setting
To support our claim that performance are comparable in an unregistered setting, we provided test results of ScanTalk when trained on the registered sequences with the dynamic chamfer distance in Table 1. In addition, in Table 2(b), we report results when trained with remeshed VOCAset meshes as detailed in Section 4.2. These results demonstrate that the unregistered setting is almost as effective as the registered one.
Table 2: Comparison of animation-to-scan distances and the effect of unregistered training losses.
(a) Distance between generated animations and scan motions. Here, $\sigma=0.1$ for the Varifold distance
HD $\downarrow$ CD $\downarrow$ Varifold $\downarrow$ ST on registered VOCA 0.031 0.00126 5.0342 ST on scans 0.020 0.00010 0.0157
(b) Quantitative results on VOCAset using different unregistered losses.
LVE $\downarrow$ MVE $\downarrow$ FDD $\downarrow$ CD (on original VOCA) 3.175 0.887 2.621 CD (on remeshed VOCA) 3.332 0.904 1.126 MSE (on original VOCA) 3.012 0.861 2.400
Finally, we highlight that “real” ground truth motions are determined by the 4D scanned data. With this in mind, in Table 2(a), we compared this raw ground truth motions with predicted motions on its corresponding registration (“ST on registered VOCA”) or directly on the first frame of the raw 4D scan sequence (“ST on scans”).
4.3.3 Ablation study
Finally, to investigate the impact of each component of the proposed model, we report the results of an ablation study in Table 3. In particular, we compare the performance with other audio encoders (Wav2Vec2 [45] and WavLM [7]), with other audio stream processing modules such as a Transformer Decoder (TD) and Bidirectional Gated Recurrent Unit (BiGRU). Finally, we report results with a more elaborated fusion strategy for the audio and geometric features using a cross-attention mechanism (CA Fusion). Notably, for this last point, we observed that this additional complexity did not lead to an uniform improvement in the quantitative results: performances are much better on VOCAset but way worse on Multiface so we decided to stick with a simple concatenation of the audio and geometric features.
Table 3: ScanTalk (ST) multi-dataset ablation studies. Top: ST with different audio encoders. Middle: ST using different audio stream processing. Bottom: ST with an other feature fusion strategy. The last row is the proposed ScanTalk

### v3

@@ 4 Experiments
In the following sections, we first outline the experimental methodology adopted for both registered and unregistered settings (Section 4.1). We then present a comprehensive quantitative evaluation of our approach in Section 4.2 and Section 4.4. We also present the results of a user study in Section 4.3. Our models are trained and evaluated on three publicly available datasets: VOCAset [VOCA2019], BIWI [BIWI_2010], and Multiface [wuu2022multiface]. Additional details on these datasets can be found in the supplementary material.
4.1 Experimental Protocols
@@
Registered setting. In this setting, all meshes of a given sequence are aligned to a common topology. This allows using a supervised loss such as the $\mathcal{L}_{MSE}$, $\mathcal{L}_{M}$, $\mathcal{L}_{Vel}$ and $\mathcal{L}_{Cos}$. Even though the topology may vary from one sequence to another, when training a single model on several registered datasets the only requirement is to rigidly align and scale the data with a reference. For our experiments, we rigidly aligned all the meshes with the VOCAset meshes.
@@
Unregistered setting. Here, we replace $\mathcal{L}_{MSE}$ with the dynamic Chamfer distance $\mathcal{L}_{CD}$. To showcase the results of a training procedure on unregistered meshes, we chose to randomly remesh VOCAset by upsampling and downsampling the meshes.In doing so, each mesh of each sequence has its own structure, different from all the others. For other settings, such as training directly on scans, one should take care about rigid alignment with the training data and the existence of a clear aperture for the mouth. Also, downsampling the training meshes can make the training faster.
4.2 Registered Training
@@
Table 1 shows the performance of ScanTalk when trained with various loss combinations across all three registered datasets in a multidataset setting. The table reports both the proposed metrics and the standard ones. We compare against state-of-the-art methods and vanilla ScanTalk trained separately on each dataset. At a first glance, it can be easily noted that none of the compared approaches outperforms the others with respect to all the metrics. In fact, while standard metrics (LVE, MVE, FDD) mostly focus on geometric measures, they do not explicitly account for the dynamics of the animation. The proposed ones instead (DTW, DFD, $\delta_{M}$, $\delta_{Cd}$) provide such information. This points towards the idea that there is still large room for improvements, but without proper measurements such assessment is difficult.
@@
Table 1: Results across different datasets. MV stands for mask+velocity, MVC for mask+velocity+cosine. State-of-the-art models are trained on a single dataset, Scantalk is trained both on single (sd) and multi-dataset (md) settings.
VOCAset BIWI6 Multiface
LVE MVE FDD DTW DFD $\delta_{M}$ $\delta_{Cd}$ LVE MVE FDD DTW DFD $\delta_{M}$ $\delta_{Cd}$ LVE MVE FDD DTW DFD $\delta_{M}$ $\delta_{Cd}$
VOCA 6.99 0.98 2.66 1.77 7.41 1.39 1.02 5.74 2.59 41.5 1.60 7.89 1.47 0.67 4.92 2.76 55.78 1.56 6.51 0.86 1.23
FaceFormer 6.12 0.93 2.16 1.33 5.39 0.86 0.58 4.08 2.16 37.1 1.56 6.61 0.87 0.55 2.45 1.45 20.24 0.87 4.13 0.22 0.73
FaceDiffuser 4.35 0.90 2.43 1.73 6.83 0.81 0.60 4.02 2.12 39.6 1.55 6.50 0.85 0.56 3.55 2.39 29.16 1.45 5.24 0.28 0.77
CodeTalker 3.55 0.89 2.26 1.33 5.66 0.80 0.55 5.19 2.64 20.6 1.50 6.61 0.91 0.59 4.09 2.38 47.90 1.44 5.95 0.55 0.96
SelfTalk 5.61 0.91 2.32 1.25 5.43 0.72 0.56 3.63 2.06 35.5 1.60 6.93 1.03 0.66 2.28 1.90 37.43 0.96 4.18 0.25 0.75
ScanTalk (md) 6.38 0.99 2.10 1.39 5.71 0.75 0.60 4.04 2.05 40.0 1.47 6.48 0.97 0.60 2.43 1.67 32.20 0.90 4.08 0.23 0.74
+MV (md) 7.03 1.09 1.47 1.14 5.16 0.78 0.58 4.75 2.22 39.4 1.51 6.74 1.09 0.63 2.54 1.89 15.97 0.99 4.21 0.27 0.77
+MVC (md) 7.05 1.07 1.29 1.19 5.25 0.72 0.52 4.44 2.09 36.6 1.46 6.71 1.01 0.61 2.37 1.71 16.00 0.93 4.71 0.33 0.74
ScanTalk (sd) 3.01 0.86 2.40 1.20 5.25 0.73 0.55 4.65 2.14 36.0 1.56 7.22 1.05 0.63 2.65 1.87 64.43 0.96 4.27 0.57 0.93
One interesting result that supports our claim is that in Table 1 vanilla ScanTalk performs better than ScanTalk trained with the additional losses (masked, velocity and cosine) in terms of geometric measures such as LVE. However, if one looks at the charts in Figure 3, which reports both visual results and the generated $y$-coordinate movements of the upper and lower lips against ground truth, can clearly observe that the dynamics of the animation is better modeled when the additional losses are used. This is reflected in lower DTW, DFD, $\delta_{M}$ and $\delta_{Cd}$ measures in Table 1, which represent a quantitative alternative to the qualitative results in the charts.
Figure 3: Comparison of ScanTalk variants and state-of-the-art methods on a VOCAset test sequence, focusing on vertical lip movements ($y$-coordinate). Each row shows visual outputs and corresponding lip-sync plots. The $y$-coordinate captures the main dynamics of lip motion. ScanTalk benefits from additional loss terms, yielding more accurate and expressive lip movement. Here same colors represent same timestep.
@@
Another notable observation from Table 1 is that while standard metrics suggest that training vanilla ScanTalk on a single dataset yields better performance, our proposed metrics reveal the opposite: the multi-dataset training setting leads to improved performance, even on individual datasets. Overall, ScanTalk trained in a multidataset setting performs favorably with respect to state-of-the-art models, mostly on VOCAset, with the newly introduced metrics providing a more complete evaluation. For example, in VOCAset, FaceDiffuser and CodeTalker report the lowest LVE and MVE. However, their motion dynamics is quite compromised when looking at the generated animations. To support this, Figure 3 reports their motion charts for a randomly picked sequence of VOCAset. The generated motion fails to a large extent in following that of the ground-truth. Indeed, this is again reflected in worse DTW, DFD and $\delta$ measures. ScanTalk trained with the additional losses, instead, better captures the motion dynamics. This improvement aligns with our expectations, as the new losses were specifically designed to enhance dynamic animation realism, a nuance that the standard metrics failed to capture. This comparison highlights the limitations of standard metrics, demonstrating their inadequate performance. In contrast, the new metrics we propose offer a significantly improved assessment of lip-sync quality. Heatmaps in Figure 4(a), which show the MSE between the ground truth and predictions on VOCAset, qualitatively reveal that ScanTalk performs comparably with other state-of-the-art methods.
In $BIWI_{6}$ and Multiface performance vary more. Specifically, ScanTalk surpasses the state-of-the-art method on $BIWI_{6}$ in DTW and DFD, but not in $\delta_{M}$ and $\delta_{Cd}$. For Multiface, ScanTalk outperforms state-of-the-art methods for the DFD metric and performs comparably in terms of other metrics. These observations align with findings from [nocentini2024scantalk3dtalkingheads], and we attribute this performance discrepancy to the significant differences among the three registered datasets. VOCAset features static faces with only mouth movements, while $BIWI_{6}$ and particularly Multiface involve notable head movements in their sequences. These dataset variations contribute to the model’s reduced performance as they introduce a further level of complexity to be modeled in a unified training scheme.
To evaluate the impact of such head rotations and the robustness of ScanTalk, we performed a specific test. We computed the relative difference $\Delta_{R}$ between a metric $M$ (e.g., LVE or others) computed on the aligned ($Original$) test set of VOCASET and when applying random rigid rotations by a certain angle ($Rotated$):
$\Delta_{R}=\frac{M(Rotated)-M(Original)}{M(Original)}.$ (22)
The results in Table 2 provide clear insights of the robustness of ScanTalk with respect to pose variation at test time. We observe that the error increase remains relatively modest even under significant rotations of the input mesh. When analyzing the LVE and MSE across rotations of up to $\pm 30^{\circ}$, the relative performance degradation is especially small for rotations around the $Y$ (yaw) and $Z$ (roll) axes.
Table 2: Increase (%) in LVE and MSE after applying random test-time rotations on VOCAset.
LVE MSE
Angle (deg.) $X$ $Y$ $Z$ $X$ $Y$ $Z$
$\pm 2^{\circ}$ 0.4 0.1 0.1 0.1 0.1 0.1
$\pm 5^{\circ}$ 1.4 0.4 0.2 0.2 0.1 0.1
$\pm 10^{\circ}$ 4.2 1.0 0.5 0.4 0.3 0.8
$\pm 20^{\circ}$ 5.4 2.2 2.0 1.5 1.0 1.6
$\pm 30^{\circ}$ 14.9 4.8 5.0 3.8 2.6 2.3
As we see in Table 2, the LVE error increases more significantly with larger rotations, especially along the $X$ (pitch) axis. The figure also shows the $XYZ$ coordinate system used to define rotation axes, overlaid on a sample VOCAset face, which helps interpret the directional impact of each rotation. This stability indicates that the model is not overfitting to rigid, canonical head poses seen during training, but rather learns representations that are invariant to moderate pose shifts. In this sense, it acts as a form of data augmentation, which improves the overall robustness of the model. More details about this experiment can be found in the supplementary material.
@@
(a) Comparison of ScanTalk with SOTA methods on VOCAset. Blue/red hues = low/high error.
@@
(b) MDS visualization of latent combinations. Same identity-audio pairs show highly similar representations, despite different parameterization. Different identities also exhibit similar MDS trajectories.
@@
Finally, to illustrate the consistency of the learned space-time feature space, in Figure 4(b) we present the Multidimensional Scaling (MDS) [torgerson1952multidimensional] projection of three sequences of latent combinations from ScanTalk, each generated with the same audio snippet but different faces to animate. The results show that when two faces share the same identity but differ in topology, their MDS sequences are very similar, nearly overlapping. In contrast, when animating different identities with the same audio, the MDS paths exhibit similar patterns but occupy distinct positions in space. These results underscore ScanTalk capability to animate faces independently of topology, focusing on the face identity. From Table 1, it is evident that ScanTalk, when trained with the Chamfer loss in a registered setting, yields results comparable to those obtained using registered losses in the same setting. This comparison underscores the robustness and adaptability of our framework.
4.3 User study
Table 3: User study on lip-sync quality: percentage of ratings per method across five categories.
VP(%) P(%) F(%) G(%) VG(%)
FaceFormer 14.9 30.6 25.5 24.5 4.5
FaceDiffuser 40.1 39.9 14.3 5.7 0.0
CodeTalker 1.2 16.3 31.4 30.2 20.9
SelfTalk 0.0 4.5 30.9 45.5 19.1
ScanTalk 1.3 10.3 26.0 38.5 23.9
With the goal of validating our claim regarding the need for additional metrics to measure the realism of animations, we designed a user-based study so to relate the subjective perception to the quantitative metrics described in Section 3.5.1 and reported in Table 1. In particular, 14 participants were shown 60 animations (all in FLAME topology, so to avoid potential biases) and asked them to evaluate the lip-sync quality and realism using a five-point scale: Very Poor (VP), Poor (P), Fair (F), Good (G), and Very Good (VG). We compared ScanTalk against 4 other state-of-the-art models [Fan_Lin_Saito_Wang_Komura_faceformer_2022, peng2023selftalk, FaceDiffuser_Stan_MIG2023, xing2023codetalker]. Note that we did not conduct the usual A/B study where users are asked which result they like more among multiple options; each animation is here given an independent score. The goal of this design is twofold: to assess to what extent (i) subjective perception correlates with the quantitative metrics, and (ii) the proposed dynamic losses affect the perceived realism with respect to other approaches.
Table 3 summarizes the user study results, showing the distribution of scores assigned to each method. Overall, the findings reveal a strong alignment between subjective perception and our proposed metrics, with ScanTalk receiving the highest proportion of Very Good ratings. Notably, FaceDiffuser, despite achieving a lower LVE than both ScanTalk and SelfTalk, received significantly lower user ratings, highlighting the inadequacy of standard metrics in capturing perceived visual quality. In contrast, the user scores show a clear correlation with our proposed metrics, underlying their effectiveness in capturing the quality of lip motion dynamics.
4.4 Unregistered Setting
Training on raw scans directly is too challenging for several reasons. Teeth, hair, occlusions, and other physical attributes can introduce significant interference in the signal, making it difficult for the model to accurately learn lip movements. While holes and noise in the raw scans are generally manageable, features like teeth or hair can be problematic. The model typically collapses or overfits unless extra information is provided through multi-modal approaches for instance. However, our approach is still a significant step forward regarding the robustness and generalizability of deep learning-based 3D talking heads. To support our claim that performance remains almost unchanged in an unregistered setting and to mimic raw scans attributes, we provide results of ScanTalk when trained on the three datasets where these undergo several types of alterations. In particular, we measure the performance when the training datasets are Registered (Rgt), Remeshed (Rmsh) by subdividing the mesh and then decimating it, randomly creating holes (H) or adding Gaussian Noise (GN) to vertex positions. More details on the data preparation are given in the supplementary material.
Table 4: Results across different datasets for the unregistered setting in different scenarios. All models are trained on the three datasets with different alteration: Registered (Rgt), Remeshed (Rmsh), randomly creating holes (H) and adding Gaussian Noise (GN) to vertex positions.